OpenStudy (anonymous):

Determine whether the given system has only one solution, no solution, or infinite # of solutions. If the solution has only one solution, enter the solution 6x-4y+4z=4 4x+2y+6z=5 5x-y+5z=-2

OpenStudy (radar):

6x-4y+4z=4 8x+4y+12z=10 (2nd equation doubled) now add getting 14x+16z=14 y has been eliminated. Now eliminate y again 4x+2y+6z=5 10x-2y+10z=-4 (last equation doubled) now add 14x +16z =1 Note that 14x+16z =1 but we have shown 14x+16=14 So it is inconsistent. I don't believe there is a solution.

OpenStudy (anonymous):

CORRECT

OpenStudy (radar):

Did you come to the same conclusion?

OpenStudy (anonymous):

yes

OpenStudy (radar):

Very good.

OpenStudy (precal):

I also, came up with the same solution with my trusty TI-84